# Unified Lorentz-covariant Poisson Bracket for the Electrodynamics of a Point Particle

### v1

@@ 3 Proposal for a Lorentz-covariant Poisson Bracket
$\{A,B\}({s})=a\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}k_{\mu}\left(\frac{\partial{A}_{s}}{\partial{q}^{\nu}(s)}\frac{\partial{B}_{s}}{\partial\pi_{\mu\nu}(s)}-\frac{\partial{B}_{s}}{\partial{q}^{\nu}(s)}\frac{\partial{A}_{s}}{\partial\pi_{\mu\nu}(s)}\right),$ (20)
where $A_{s}$ and $B_{s}$ denote that functions $A$ and $B$ have been written in terms of the initial conditions $A(s)$ and $B(s)$, respectively, with $s$ some fixed instant. In the present case, and for the free field, we have that $A_{\mu}$ is a linear function of the initial conditions. Consequently, it is straightforward to calculate the non-equal-time pb for $A_{\mu}$ at two different space-time positions, so that we obtain that
$\{A_{\mu}(x_{a}),A_{\nu}(x_{a}^{\prime})\}({s})=a\pi{c}\left[\frac{\delta^{2}(0)}{8\pi^{3}}\right]\eta_{\mu\nu}\Delta(x_{\alpha}-x_{\alpha}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part},$ (21)
where $\Delta$ denotes the Lorentz-invariant Pauli-Jordan function. Similarly, one can directly calculate the non-equal-time pb of the vector potential and its conjugate momentum:
$\{A_{\mu}(x_{a}),\theta_{\lambda\nu}(x_{a}^{\prime})\}({s})=\frac{a}{4}\left[\frac{\delta^{2}(0)}{8\pi^{3}}\right]\eta_{\mu\nu}\partial_{\lambda}\Delta(x_{\alpha}-x_{\alpha}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part}.$ (22)
Both results are independent of $s$ and coincide, except for a numerical factor [^5] The factor $\delta^{2}(0)/8\pi^{3}$ in Eqs. (17), (19), (21), and (22), or similar factors, will appear when calculating pbs with respect to the canonical variables of quantities consisting of sums (integrals) of the contributions of all the modes of oscillation of the field, and is due to the particular choice of the normalization factor for these quantities. This factor could be “eliminated” through a renormalization process, but instead of sweeping it under the rug, we have decided to leave it explicit., with their counterparts from the conventional formulations of field theories (see, for example, Heitler [21]).
One more comment on Eqs. (21) and (22) is worth making as they show that the proposal in Eq. (20) leads to manifestly Lorentz-covariant results. Thus, although the proportionality of $V_{\mu}$ to $k_{0}$ might seem to break the covariance of the expression, we have actually that such proportionality transforms the pb with respect to variables $q_{\mu}$ and $\pi_{\mu\nu}$ into an effectively Lorentz-covariant form.
4 Reduction of the Proposal to the Standard (Non-covariant) Form of the Poisson Bracket
With the aim of finding an appropriate value for the proportionality constant $a$, let us explore the reduction process of the pb in Eq. (15) to its non-covariant form. First, and for purposes that will be clarified later, we rewrite Eq. (15) in terms of amplitudes $\mathcal{A}_{\mu}$ and their complex conjugates, so that we establish that
$\{A,B\}_{q\pi}=4\pi{i}c\{A,B\}_{\mathcal{AA}^{\ast}},$ (23)
@@ 4 Reduction of the Proposal to the Standard (Non-covariant) Form of the Poisson Bracket
$\{A,B\}_{\mathcal{AA}^{\ast}}=a\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}^{2}\left(\frac{\partial{A}}{\partial\mathcal{A}^{\mu}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\mu}}-\frac{\partial{B}}{\partial\mathcal{A}^{\mu}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\mu}}\right).$ (24)
Naturally, Eq. (24) does not correspond to the pb associated with the normal variables of the electromagnetic field, since for each oscillation mode of the field there is a sum of four terms instead of two, as occurs in the non-covariant formulation.
@@
$\{A,B\}_{\mathcal{AA}^{\ast}}=a\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}^{2}\left(\sum_{i=1}^{3}\left[\frac{\partial{A}}{\partial\mathcal{A}_{i}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{i}}-\frac{\partial{B}}{\partial\mathcal{A}_{i}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{i}}\right]-\left[\frac{\partial{A}}{\partial\mathcal{A}_{0}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{0}}-\frac{\partial{B}}{\partial\mathcal{A}_{0}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{0}}\right]\right).$ (26)
Then, since the Lorentz auxiliary condition, $\partial_{\mu}A^{\mu}=0$, implies that
@@
we have that the contributions of the scalar modes to the pb are compensated by those of the longitudinal modes for each oscillation mode of the field. Therefore, we find through this reduction process (Gupta-Bleuler process [22, 23]) that the pb in Eq. (26) is equal to
$\{A,B\}_{\mathcal{AA}^{\ast}}=\frac{a}{2}\sum_{\lambda=1}^{2}\int{d}^{3}k\hskip 1.42262ptk_{0}\left(\frac{\partial{A}}{\partial\mathcal{A}_{\lambda}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\lambda}}-\frac{\partial{B}}{\partial\mathcal{A}_{\lambda}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\lambda}}\right),$ (28)
with $k_{0}=|\mathbf{k}|$.
Now, the pb in Eq. (28) may not look like the standard pb, but it is in fact equal to the latter. To show this equivalence between expressions, we need to remember that the amplitudes in Eq. (28) are still four-dimensional and that the relation of these amplitudes with their three-dimensional counterparts is precisely [^6] See, for example, Heitler [21] or Bogoliubov and Shirkov [24].
@@
so that we finally obtain
$\{A,B\}_{\mathcal{AA}^{\ast}}=\frac{a}{4}\sum_{\lambda=1}^{2}\int{d}^{3}k\left(\frac{\partial{A}}{\partial\mathcal{A}_{\lambda}(\mathbf{k})}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\lambda}(\mathbf{k})}-\frac{\partial{B}}{\partial\mathcal{A}_{\lambda}(\mathbf{k})}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\lambda}(\mathbf{k})}\right),$ (30)
which is the pb associated with the normal variables of the electromagnetic field if we take $a=4$ and identify $\lambda$ as the polarization index.
It is very important to note that although our proposal converges correctly to the standard pb, we have, due to Eq. (29), that
@@
$\{\mathcal{A}_{\mu}(k_{\alpha}),\mathcal{A}^{\ast}_{\nu}(k_{\alpha}^{\prime})\}_{\mathcal{AA}^{\ast}}=4\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}^{2}\eta_{\mu\nu}\delta^{(4)}(k_{\alpha}-k_{\alpha}^{\prime}).$ (32)
The pb in the right-hand side of Eq. (31) is generally used as the starting point for canonical quantization in the conventional formulations of quantum field theories. This primacy is the reason why the reduction process of the proposed pb was done in terms of amplitudes rather than canonical variables. It is particularly interesting to note that our result for the pb of two amplitudes of the electromagnetic field, Eq. (32), coincides in structure with that obtained by Dirac through a completely different approach [25].
@@ 5 Joint Description of the Electrodynamics of a Point Particle
$\{A,B\}_{xp}=\frac{\partial{A}}{\partial{x}_{\mu}}\frac{\partial{B}}{\partial{p}^{\mu}}-\frac{\partial{A}}{\partial{x}_{\mu}}\frac{\partial{B}}{\partial{p}^{\mu}},$ (33)
@@
$\{x_{\mu},p_{\nu}\}_{xp}=\eta_{\mu\nu}.$ (34)
On this basis, we establish that the set $(Q;P)$ of canonical variables of the particle+electromagnetic field system should be the union of the sets of canonical variables of the mechanical subsystem and the field, that is,
$(Q;P)=(x_{\mu},q_{\mu}(k_{\alpha});p_{\mu},\pi_{\mu\nu}(k_{\alpha})).$ (35)
@@
$\{A,B\}_{QP}(s)=\{A,B\}_{xp}(s)+\{A,B\}_{q\pi}(s)$ (36)
as the pb associated with all the canonical variables at time $s$ and postulate such bracket as the basis for a joint description of the whole system that is both canonical and Lorentz-covariant.
An interesting feature of the proposed description comes from the fact that the canonical variables of the system at any two different times, namely $s$ and $\tau>s$, are related through a canonical transformation. Thus, we have that
$\{x_{\mu}(\tau),p_{\nu}(\tau)\}_{QP}(s)=\{x_{\mu}(\tau),p_{\nu}(\tau)\}_{QP}(\tau),$ (37)
since pbs are invariant under those transformations. Therefore, we straightforwardly establish from Eqs. (34) and (37) that
$\{x_{\mu}(\tau),p_{\nu}(\tau)\}_{xp}(s)+\{x_{\mu}(\tau),p_{\nu}(\tau)\}_{q\pi}(s)=\eta_{\mu\nu},$ (38)
i.e, that the dependence of the canonical variables of the mechanical subsystem at time $\tau$ (or at any other time) on the initial conditions $x_{\mu}(s)$, $p_{\mu}(s)$, $q_{\mu}(s)$, and $\pi_{\mu\nu}(s)$ is precisely of the form necessary to ensure that their pb equals the metric tensor.
@@ 6 Conclusions
The apparent incompatibility between the requirement of field theories for locally describing the interactions between particles and the extended use of the canonical quantization process seems to have found, at least in principle, a solution in the mhf. As we showed in this article, it is feasible to use this formalism to construct a description of the complete (particle++electromagnetic field) system that is suitable for a canonical treatment and, at the same time, treats all space-time coordinates symmetrically. The pivotal element of such a description is an unambiguous Lorentz-covariant pb, Eqs. (15) and (20), which not only reproduces the well-known results from the conventional formulations of field theories, but also converges to the standard pb when reduced to its non-covariant form. On this basis, we obtained the pb of the canonical variables of the particle and of the field, Eqs. (16) and (34). These results constitute a generalization of the canonical pb of non-relativistic classical mechanics and might be taken as the departure point for the future development of a manifestly Lorentz-covariant quantization process.

### v2

@@ 3 Proposal for a Lorentz-covariant Poisson Bracket
$\{A,B\}({s})=a\int{d}^{4}k\hskip 1.42262ptk_{\mu}\left(\frac{\partial{A}_{s}}{\partial{q}^{\nu}(s)}\frac{\partial{B}_{s}}{\partial\pi_{\mu\nu}(s)}-\frac{\partial{B}_{s}}{\partial{q}^{\nu}(s)}\frac{\partial{A}_{s}}{\partial\pi_{\mu\nu}(s)}\right),$ (23)
where $A_{s}$ and $B_{s}$ denote that functions $A$ and $B$ have been written in terms of the initial conditions $A(s)$ and $B(s)$, respectively, with $s$ some fixed instant. In the present case, and for the free field, we have that $A_{\mu}$ is a linear function of the initial conditions; therefore, it is straightforward to calculate the non-equal-time PB for $A_{\mu}$ at two different space-time positions, so that we obtain that
$\{A_{\mu}(x_{a}),A_{\nu}(x_{a}^{\prime})\}({s})=2\pi{a}{c}\left[\frac{\delta(0)}{8\pi^{3}}\right]\eta_{\mu\nu}\Delta(x_{\alpha}-x_{\alpha}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part},$ (24)
where $\Delta$ denotes the Lorentz-invariant Pauli–Jordan function. Similarly, one can directly calculate the non-equal-time PB of the vector potential and its conjugate momentum,
$\{A_{\mu}(x_{a}),\theta_{\lambda\nu}(x_{a}^{\prime})\}({s})=\frac{a}{2}\left[\frac{\delta(0)}{8\pi^{3}}\right]\eta_{\mu\nu}\partial_{\lambda}\Delta(x_{\alpha}-x_{\alpha}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part}.$ (25)
Both results are independent of $s$ and coincide, except for a numerical factor [^4] The factor $\delta(0)/8\pi^{3}$ in Eqs. (20), (22), (24), and (25), or similar factors, will appear when calculating PB with respect to the canonical variables of quantities consisting of sums (integrals) of the contributions of all the modes of oscillation of the field, and is due to the particular choice of the normalization factor for these quantities. This factor could be “eliminated” through a renormalization process, but instead of sweeping it under the rug, we have decided to leave it explicit., with their counterparts from the conventional formulations of field theories (see, for example, Heitler [33]).
4 Reduction of the Proposal to Its Non-covariant Form
With the aim of finding an appropriate value for the proportionality constant $a$, let us explore the reduction process of the PB in Eq. (18) to its non-covariant form. First, and for purposes that will be clarified later, we rewrite Eq. (18) in terms of the amplitudes $\mathcal{A}_{\mu}$ and their complex conjugates, so that we straightforwardly establish that
$\{A,B\}_{q\pi}=-4\pi{i}c\{A,B\}_{\mathcal{AA}^{\ast}},$ (26)
@@ 4 Reduction of the Proposal to Its Non-covariant Form
$\{A,B\}_{\mathcal{AA}^{\ast}}=a\int{d}^{4}k\hskip 1.42262ptk_{0}\left(\frac{\partial{A}}{\partial\mathcal{A}^{\mu}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\mu}}-\frac{\partial{B}}{\partial\mathcal{A}^{\mu}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\mu}}\right).$ (27)
Naturally, Eq. (27) does not correspond to the standard (non-covariant) PB associated with the normal variables of an electromagnetic radiation field, since for each oscillation mode of the field there is a sum of four terms instead of two.
@@
$\{A,B\}_{\mathcal{AA}^{\ast}}=a\int{d}^{4}k\hskip 1.42262ptk_{0}\left(\frac{\partial{A}}{\partial\mathcal{A}_{0}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{0}}-\frac{\partial{B}}{\partial\mathcal{A}_{0}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{0}}-\sum_{i=1}^{3}\left[\frac{\partial{A}}{\partial\mathcal{A}_{i}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{i}}-\frac{\partial{B}}{\partial\mathcal{A}_{i}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{i}}\right]\right).$ (29)
Then, since we are working in the Lorentz gauge, we have that the Lorentz auxiliary condition, $\partial_{\mu}A^{\mu}=0$, implies that
@@
so the contributions of the scalar modes to the PB are compensated by those of the longitudinal modes for each oscillation mode of the field. Therefore, we find through this reduction process (Gupta–Bleuler process [34, 35]) that the PB in Eq. (29) is equal to
$\{A,B\}_{\mathcal{AA}^{\ast}}=-a\sum_{\lambda=1}^{2}\int{d}^{4}k\hskip 1.42262ptk_{0}\left(\frac{\partial{A}}{\partial\mathcal{A}_{\lambda}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\lambda}}-\frac{\partial{B}}{\partial\mathcal{A}_{\lambda}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\lambda}}\right).$ (31)
Now, the PB in Eq. (31) is certainly not the standard PB for an electromagnetic radiation field, but both expressions are in fact related. To show this, we need to remember that the amplitudes in Eq. (31) are still four-dimensional and that the relation of these amplitudes with their three-dimensional counterparts is precisely [^5] See, for example, Heitler [33] or Bogoliubov and Shirkov [36].
@@
so that
$\{A,B\}_{\mathcal{AA}^{\ast}}=-\frac{a}{2}\sum_{\lambda=1}^{2}\int{d}^{4}k\left(\frac{\partial{A}}{\partial\mathcal{A}_{\lambda}(\mathbf{k})}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\lambda}(\mathbf{k})}-\frac{\partial{B}}{\partial\mathcal{A}_{\lambda}(\mathbf{k})}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\lambda}(\mathbf{k})}\right).$ (33)
From this result, we immediately establish that
$\{A,B\}_{\mathcal{AA}^{\ast}}=-\frac{a}{2}\int{d}k_{0}\left[\sum_{\lambda=1}^{2}\int{d}^{3}k\hskip 1.42262pt\left(\frac{\partial{A}}{\partial\mathcal{A}_{\lambda}(\mathbf{k})}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\lambda}(\mathbf{k})}-\frac{\partial{B}}{\partial\mathcal{A}_{\lambda}(\mathbf{k})}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\lambda}(\mathbf{k})}\right)\right].$ (34)
Thus, we find that our proposal is equal to the integral over $k_{0}$ of the standard PB for an electromagnetic radiation field if we take a=2a=-2 and identify $\lambda$ as the polarization index. A relationship involving an integral is to be expected since the standard PB involves an integral over the field’s three-dimensional wave vector, which is not Lorentz-covariant, so this extra operation guarantees the covariance of our proposal.
It is very important to note that although our proposal indeed contains the standard PB for an electromagnetic radiation field, we have, due to Eq. (32), that
@@
$\{\mathcal{A}_{\mu}(k_{\alpha}),\mathcal{A}^{\ast}_{\nu}(k_{\alpha}^{\prime})\}_{\mathcal{AA}^{\ast}}=-2k_{0}\eta_{\mu\nu}\delta^{(4)}(k_{\alpha}-k_{\alpha}^{\prime}).$ (36)
The PB in the right-hand side of Eq. (35) is generally used as the starting point for canonical quantization in the conventional formulations of quantum electrodynamics. This primacy is the reason why the reduction process of the proposed PB was done in terms of amplitudes rather than canonical variables. It is particularly interesting to note that our result for the PB of two amplitudes of the electromagnetic radiation field, Eq. (36), coincides in structure with that obtained by Dirac through a completely different approach [37].
@@ 5 Joint Description of the Electrodynamics of a Point Particle
$\{A,B\}_{up}=-\left(\frac{\partial{A}}{\partial{u}_{\mu}}\frac{\partial{B}}{\partial{p}^{\mu}}-\frac{\partial{B}}{\partial{u}_{\mu}}\frac{\partial{A}}{\partial{p}^{\mu}}\right),$ (37)
@@
$\{u_{\mu},p_{\nu}\}_{up}=-\eta_{\mu\nu}.$ (38)
On this basis, we establish that the set $(Q;P)$ of canonical variables of the point particle+electromagnetic radiation field system should be the union of the sets of canonical variables of the mechanical subsystem and the field, that is,
$(Q;P)=(u_{\mu},q_{\mu}(k_{\alpha});p_{\mu},\pi_{\mu\nu}(k_{\alpha})).$ (39)
@@
$\{A,B\}_{QP}(s)=\{A,B\}_{up}(s)+\{A,B\}_{q\pi}(s)$ (40)
as the PB associated with all the canonical variables at instant $s$ and postulate such bracket as the basis for a joint description of the complete system that is both canonical and Lorentz-covariant.
An interesting feature of the proposed description comes from the fact that the canonical variables of the system at any two different instants, namely $s$ and $\tau>s$, are related through a canonical transformation. Thus, we have that
$\{u_{\mu}(\tau),p_{\nu}(\tau)\}_{QP}(s)=\{u_{\mu}(\tau),p_{\nu}(\tau)\}_{QP}(\tau),$ (41)
since PB are invariant under those transformations. Therefore, we straightforwardly establish from Eqs. (38) and (41) that
$\{u_{\mu}(\tau),p_{\nu}(\tau)\}_{up}(s)+\{u_{\mu}(\tau),p_{\nu}(\tau)\}_{q\pi}(s)=-\eta_{\mu\nu},$ (42)
i.e, that the dependence of the canonical variables of the mechanical subsystem at instant $\tau$ (or at any other instant) on the initial conditions $u_{\mu}(s)$, $p_{\mu}(s)$, $q_{\mu}(s)$, and $\pi_{\mu\nu}(s)$ is precisely of the form necessary to ensure that their PB equals the metric tensor.
@@ 6 Conclusions
The apparent incompatibility between the requirement of field theories for locally describing the interactions between particles and the extended use of the canonical quantization process seems to have found, at least in principle, a solution in the MHF. As we showed in this article, it is feasible to use this formalism to construct a description of the complete (particle++electromagnetic radiation field) system that is suitable for a canonical treatment and, at the same time, treats all space-time coordinates symmetrically. The pivotal element of such a description is an unambiguous Lorentz-covariant PB, Eqs. (18) and (23), which not only reproduces the well-known results from the conventional formulations of field theories, but also contains the standard (non-covariant) PB for an electromagnetic radiation field when reduced to its non-covariant form. On this basis, we obtained the PB of the canonical variables of the particle and of the field, Eqs. (19) and (38). These results constitute a generalization of the canonical PB of non-relativistic classical mechanics and might be taken as the departure point to generalize Cetto’s approach to the quantization of the electromagnetic radiation field and, therefore, develop a manifestly Lorentz-covariant quantization process for such a field.
The generalization of the PB in Eqs. (18) and (23) for arbitrary values of the gauge-fixing parameter, other than that associated with the Lorentz gauge, remains to be done in order to construct a Lorentz-covariant generalization of canonical quantization for general fields.
Acknowledgments: The author would like to thank two anonymous reviewers for their very helpful comments.
Funding information: Not applicable.